(x^2 - 1)dy - (y^2 - 1)dx


1
Expert's answer
2021-07-09T12:53:02-0400

Solving the equation

"(x^2 - 1)dy - (y^2 - 1)dx=0"

We can rewrite as:


"(x^2-1)dy=(y^2-1)dx"

Dividing by "(x^2-1)(y^2-1)" through:


"\frac{d y}{y^2-1}=\frac{dx}{x^2-1}"

Integrate both sides :


"\int \frac{1}{y^2-1} d y=\int \frac{1}{x^2-1} dx"

Evaluate the integrals:


"\frac{1}{2} \log (-y+1)-\frac{1}{2} \log (y+1)=
\frac{1}{2} \log (-x+1)-\frac{1}{2} \log (x+1)+c_{1}"

where "c_{1}" is an arbitrary constant.


Solving for y=y(x) we have


"y=-\frac{x+e^{2 c_{1}}(x-1)+1}{-x+e^{2 c_{1}}(x-1)-1}"

Simplifying the arbitrary constant, we have


"y=-\frac{x+c_{1}(x-1)+1}{-x+c_{1}(x-1)-1}"

which is the required solution to the given differential equation.